OpenStudy (anonymous):

Find the average rate of change for the function over the given interval. y=4x^2 between x=0 and x = 7/4

OpenStudy (anonymous):

if f(x) = 4x^2, then average rate of change is given by: [ f(7/4) - f(0) ]/ (7/4 - 0)

OpenStudy (anonymous):

how do you do it with fractions im so lost

OpenStudy (anonymous):

f(7/4) = 4*(7/4)^2 = 4 * (7/4) * (7/4) = 49/4 can you give me f(0) = ?

OpenStudy (anonymous):

wait you dont multiply 4 times 4? u only square the numerator?

OpenStudy (anonymous):

that exponent is only for the 7/4...